# High-dimensional inference of radio interferometer beam patterns I: Parametric model of the HERA beams

### v1

@@ 1 Introduction
To avoid these pitfalls, our approach is as follows. Rather than expecting an unrealistically precise measurement of either the beam pattern or a reference sky model, we cast beam estimation as a Bayesian inference problem involving an explicit forward model for the observed data. The uncertainty on each component of the data model – the beam itself, the sky model, and even the direction-independent gains – can be quantified and folded into the analysis in a statistically principled way. As long as the parametric models for each component have sufficient flexibility, biases can be avoided by marginalising over the various sources of uncertainty. While this typically results in a substantially increased uncertainty in the recovered beam model, these uncertainties have the virtue of accurately reflecting the state of our knowledge about the whole system, and so highlight where modelling/measurement improvements are most needed. We present this work across two papers. In this paper, we focus on the model-building aspect of the inference, and in the second paper we demonstrate the feasibility of the inference on a toy problem.
Bayesian methods typically require an analytic parametrization of the beam, which is an active area of research in the HERA collaboration.111 http://reionization.org/manual_uploads/HERA101_Analytic_polarized_beam.pdf [^2] https://reionization.org/manual_uploads/HERA114_MEMO_beam_harmonics_Cynthia.pdf Analytic parametrizations have demonstrated significant promise in solving problems in HDR applications, e.g. more rapid beam evaluation (Asad et al., 2021), better physical beam characterization (Sekhar et al., 2022; Nasirudin et al., 2022), and removal of calibration biases that stem from gridding errors (Barry & Chokshi, 2022). This last reference also highlights the importance of accurate beam models for the sake of accurate gridding kernels in imaging power spectrum analyses (Morales et al., 2019). Bayesian instrumental characterization methods have been demonstrated before (e.g. Lochner et al., 2015; Anstey, 2023; Cumner et al., 2023). Alternatives to this one are being explored within the HERA collaboration as well (Kern et al. in prep.). Our approach has similarities with several of them, but differs on a technical level in an important way. In particular, we use a Gibbs sampling approach to perform the inference. If the models of the various components of the data are constructed carefully, this allows us to perform inference on extremely large numbers of parameters, allowing highly detailed parametric models of each individual antenna beam to be inferred while simultaneously modelling every other antenna and an uncertain sky model. This is made computationally tractable by exploiting the properties of multivariate Gaussian subspaces of the full joint posterior distribution, as we will demonstrate in the accompanying paper. While still a computationally-intensive process, this makes it possible to simultaneously model highly complicated beam structures of many individual antennas in large interferometric arrays. We implement this Bayesian modeling within the Hydra Gibbs sampling framework. [^3] https://github.com/HydraRadio/Hydra
@@ 2 Mathematical Formalism
In order to incorporate the beam into our Bayesian inference, we will need to posit a parameterized model for the beam that relates to the data from which the inference is drawn. In the general landscape of Bayesian inference, this is a rather arbitrary decision moderated only by the fact that some parametrizations may describe the data more efficiently than others. Without extreme prior knowledge about the beam, and in particular, uncertainties in the beam due to physical variations from the ideal scenario, we are left to choose some arbitrary basis for the beam with the hopes that our existing prior knowledge can at least constrain which basis functions are the most dominant. As a concrete example, we can take a detailed electromagnetic simulation or highly precise holographic measurement and fit a linear basis to the beam pattern for a high number of modes. Depending on the beam pattern and choice of basis functions, we may find that most of the fit is dominated by a small subset of basis functions. This has been shown to be relatively effective in de Lera Acedo et al. (2013); Bui-Van et al. (2017); Iheanetu et al. (2019); Asad et al. (2021), though the choice of basis is often idiosyncratic to the telescope. We can then restrict ourselves to these dominant basis functions and consider the coefficients as the free parameters in our model.
@@
Figure 1: All Jones components of the simulated Phase I dipole at 150 MHz, separated into real and imaginary components. There is an obvious dipolar structure, and higher azimuthal modes are clearly visible at zenith angles of 60 degrees or greater.
@@
An appealing alternative, though one that seems to us extremely computationally demanding for a full Bayesian computation, would be to have not only detailed knowledge of the ideal receiving element, but also have a parameterization for physical perturbations to this ideal along with detailed simulations or calculations of the response to them. Then a few parameters describing the perturbations only need to be constrained. The primary issue is that perturbations tend to belong to a multidimensional continuum and developing simulation outputs for finely graded perturbations to understand the resulting effect on the beam pattern is computationally prohibitive, though not infeasible depending on the study design (Kim et al., 2022, 2023). If a suitable emulator were constructed (i.e. if we could apply machine learning to produce fast but necessarily incomplete outputs of the simulator), the computational overhead could be reduced in the long term with an initial investment. However this might violate intellectual property law and is a somewhat ill-defined problem for a general electromagnetic simulator. Due to these seemingly intense challenges, and since our goal is mainly to establish the tractability and usefulness of our Gibbs sampling method, we opt for the linear basis expansion described in the previous paragraph.
@@
Figure 2: All Jones components of simulated HERA vivaldi feed at 150 MHz. Spatial structure is somewhat more complex than the Phase I dipoles.
@@
We aim to infer the beam parameters from the visibilities via a forward model. We write the visibility equation in terms of the polarized beam response for antenna $j$ (Jones matrix), $\boldsymbol{\mathcal{J}}_{j}(\hat{s},\nu)$, sky coherency matrix, $\boldsymbol{\mathcal{C}}(\hat{s},\nu)$, and physical antenna separation for antennas $j$ and $k$, $\Delta\vec{x}_{jk}$, as,
@@
where $c$ is the speed of light, $\nu$ is the frequency of observation, $\hat{s}$ is a unit vector pointing to different positions on the sky, $i^{2}=-1$, and depending on the coordinate system, some of these quantities are time-dependent. Different choices of coordinate system present different trade-offs. The Bayesian inference for the beam is simpler if we operate in a coordinate system where the beam pattern is fixed. For the sky coordinates, we choose azimuth, $\phi$ and zenith angle, $\theta$. We choose sky polarization directions that align with this coordinate system. This puts all of the time dependence in the sky coherency matrix at the cost of producing a discontinuity in the beam pattern at zenith. This discontinuity arises because in this coordinate system, there is no well-defined direction for the polarization unit vectors at zenith. In math, if we think of the polarization unit vectors, $\hat{\theta}$ and $\hat{\phi}$ as vector fields, then
@@
In other words the projection is area-preserving (up to an overall factor of $2\alpha$). For $\alpha=1$, $\rho=0$ maps to zenith and $\rho=1$ maps to the horizon. In this case, the total volume contributed by a basis function to the square of a Jones element is simply given by
$V_{j\hat{\alpha}nm}^{p}=\frac{|a_{j\hat{\alpha}nm}^{p}|^{2}}{2\alpha}.$ (8)
@@ 3 Compressing the Basis
There are two critical concerns when choosing a basis to represent the beam. First, assuming a priori knowledge of the beam through electromagnetic physics and knowledge of the receiving system, does the basis compactly represent the ideal beam assuming the perfect modeling of the receiving system? Second, assuming a class of perturbations to the receiving system, are ensuing perturbations to the beam pattern also compactly represented by the chosen basis? In this work, we provide a basis that satisfies the first question for the HERA receiving systems (both the Phase I dipole and Vivaldi feeds) using simulations of the ideal receiving system constructed in Fagnoni et al. (2021b, a). We leave a full treatment of the second question for future work.
@@
In order to compress the basis, we need a way of determining which modes are the most important to use. One way to quantify this is to examine equation 8 (the FBES) for the very precise fit using thousands of coefficients, shown in Figure 5 for both feed types and sky polarizations at 150 MHz. Almost all of the energy is at low $n$ and $m$ for both feeds, suggesting that we can choose a much smaller basis set with relatively little error. Since all of the plotted Jones elements have similar volume, we can interpret the number of high-energy modes as a proxy for the spatial complexity of the beam i.e. how compressible it is. For both feeds, we see that the response to azimuthally polarized radiation has longer tails to high radial modes compared to the response to zenith angle polarized radiation. Additionally, the Vivaldi feed has more higher energy modes than the Phase I dipole. For example the $|m|=3$ modes are prominent out to fairly high radial power, and there are more significant modes at much higher azimuthal number.
@@
Figure 10: (Top) Delay power spectra ($y$-axis) for the simulations described in section 4 using the HERA Dipole feed as a function of delay ($\tau$, $x$-axis) for the XX (left) and YY (right) polarizations. The solid black lines show the delay power spectra of the visibilities simulated with the CST beam from Fagnoni et al. (2021b). The colored lines show the delay power spectra for the simulations using our fits to the CST beam using various numbers of basis functions, as indicated in the colorbar. All power spectra have been normalized relative to the value of the black line at $\tau=0$. The grey shaded region marks the delays less than or equal to the Nyquist delay (500 ns) corresponding to the spectral resolution of the raw Fagnoni beam simulations (1 MHz). (Bottom) Fractional difference between the Fagnoni beam and FB simulations for each value of $N_{\text{coeff}}$. While the delay power spectra shown here correspond to a single 14.6 m EW baseline, the results are representative of all simulated baseline lengths and orientations.
@@
Figure 7 shows the RMSE as a function of zenith angle for both feed types and polarizations and a range of sparsity. We find that with 32 modes, there is $\sim 1\%$ agreement in the main lobe and $\sim 10\%$ agreement in the sidelobes. With 512 modes we achieve better than 1% precision at almost all zenith angles. The residuals from some of the sparse fits are shown in Figures 8 and 9 for the Phase I dipole and Vivaldi feeds, respectively. For a given level of sparsity, errors generally range by one order of magnitude above and below the RMSE value depending on the location of interest. We explore computational complexity for our Gibbs sampling pipeline in paper II more thoroughly, however we expect having a beam model with 64-128 parameters is computationally feasible, while 512 modes will be tractable with a more optimized implementation. We note that translating the observed errors in this fit to the actual errors produced by this approach in practice is not necessarily straightforward. The actual beam will at best be some perturbation of the simulated beam used in this work, not even including the close-packed array effects. Perhaps more importantly, the approach in paper II involves inferring the beam based on the interferometric visibilities, whereas this least-squares fit is more analogous to a problem where one infers the beam based on a holographic map. In other words, this work verifies that the spatial structure of the beam can be captured by this basis with relatively few basis functions, but the errors shown do not necessarily represent the types of errors we may observe with a Bayesian point estimate since the information content of the visibilities will be different than this simulation. Furthermore, our approach will return a full Bayesian posterior, not just a point estimate. Folding uncertainties into a forward modeling procedure should hopefully alleviate some of the difficulties that come with point estimation such as calibration errors Byrne et al. (2021).
@@ 4 Spectral Structure
We investigate any spurious spectral contamination introduced by fitting the beam with the chosen basis. To quantify this contamination, we compared the delay power spectra (Parsons & Backer, 2009; Parsons et al., 2012) from a suite of visibility simulations which use differing numbers of fit coefficients for the beam. All other input parameters to the simulations were kept fixed, however, to isolate the effects of the changing beam.
@@
The simulations were performed using pyuvsim [^5] https://github.com/RadioAstronomySoftwareGroup/pyuvsim (Lanman et al., 2019), a high-precision visibility simulator which has been rigorously tested (Lanman et al., 2022; Aguirre et al., 2022). We simulated a selection of baselines from a close-packed hexagonal array layout with various orientations and lengths ($0\leq b\leq 90$ m). For the sky, we used the 2016 Global Sky Model (GSM, Zheng et al. (2017)), a map of galactic diffuse emission. For each antenna feed type (dipole or vivaldi), we ran simulations using varied numbers of beam fit coefficients, from 32 to 512 (“FB simulations”). For a given number of fit coefficients ($N_{\text{coeff}}$), we use the FBES to select the $N_{\text{coeff}}$ modes with the largest FBES amplitudes. We do this independently at each frequency, and take the union of all basis functions selected in this way across the simulated band. Since the top $N_{\text{coeff}}$ contributing basis functions varies slightly as a function of frequency, this means each selected $N_{\text{coeff}}$ is less than the number of modes actually used in each instance of the test. Note that this is different in previous sections, where we were only examining one frequency. As a reference against which we can compare the FB simulations, we also ran a simulation for each antenna feed type using the corresponding raw electromagnetic simulations from Fagnoni et al. (2021b, a) (“Fagnoni simulations”). The delay spectra for all of these visibility simulations are shown in Figures 10 and 11 for the dipole and vivaldi feeds, respectively.
The raw beams from Fagnoni et al. (2021b, a) were simulated using a spectral resolution of 1 MHz. Our visibility simulations, however, were performed at $\sim 97$ kHz resolution to match the HERA instrument. We were thus required to interpolate the Fagnoni beams along the frequency axis to obtain the beam at the frequencies in our visibility simulations. The Nyquist frequency for a 1 MHz bandwidth is 500 ns. Delays in the range $|\tau|>500$ ns are therefore susceptible to interpolation artifacts. It is therefore plausible that any excess power we see beyond 500 ns in the delay power spectrum of the Fagnoni simulations relative to the Fourier-Bessel simulations is a result of the frequency interpolation of the Fagnoni beams. We can see potential evidence of this in Figures 10 and 11. For both feed types, there are notable spikes in the fractional difference for all values of $N_{\rm{coeff}}$ at 500 ns and 1500 ns (a harmonic of the 500 ns Nyquist frequency).
@@
It is important to note that both the raw Fagnoni beams and our FB fit coefficients were required to be interpolated in frequency for our visibility simulations. Interpolating a set of coefficients for a set of analytic functions is in principle quite different from interpolating each pixel individually in the Fagnoni beam simulations, as is done inside pyuvsim. For the results shown in Figures 10 and 11, we used cubic frequency interpolation of the FB fit coefficients. We similarly used cubic interpolation for the raw Fagnoni beams when running pyuvsim. As an additional check, we also ran a pyuvsim simulation using linear frequency interpolation of the Fagnoni beam to assess the importance of the degree of the polynomial used in the interpolation. For linear interpolation, interpolation artifacts were highly visible as a set of peaks (duplicates of the peak at $|\tau|\leq 500$ ns) spaced uniformly every 500 ns (the Nyquist frequency for 1 MHz spectral resolution). While the amplitude of these duplicated peaks decreased with increasing $|\tau|$, the dynamic range between $\tau=0$ ns and the highest delays was only $\sim 10^{8}$. Whereas for the cubically interpolated Fagnoni simulations, we see a dynamic range of $\sim 10^{11}$ in Figures 10 and 11 and little evidence of such duplicated low-delay structure.
@@
There are, however, also differences between the Fagnoni beam and FB beam simulations inside the Nyquist delay. The raw Fagnoni beams were simulated on a rectilinear grid in altitude and azimuth with a 1 deg resolution. These beams thus also require spatial interpolation during visibility simulation to get the beam value at the location of each source (pixel) in the sky catalog. Our FB beam fits are analytic and thus require no spatial interpolation. To determine the importance of this spatial interpolation on our results, we also performed a set of simulations using a mock, gridded point source catalog with one point source at the center of each Fagnoni beam pixel at a single LST. In this case, the Fagnoni beams require no spatial interpolation. For this gridded point source catalog, we see almost no discrepancy inside the Nyquist delay. This suggests that the discrepancies we see for $|\tau|\leq 500$ ns are due to the spatial interpolation of the Fagnoni beam during visibility simulation.

### v2

@@ 1 Introduction
In this paper series, we describe and demonstrate a parametric Bayesian method for inferring the per-element beam from observed visibilities. Bayesian instrumental characterization methods have been demonstrated before in the context of radio astronomy (e.g. Yatawatta, 2018; Lochner et al., 2015; Sims et al., 2022a, b; Anstey, 2023; Cumner et al., 2023; Sims et al., 2023). In this paper (Paper I), we focus on the model-building aspect of the inference, and in the second paper (Paper II) we demonstrate the feasibility of the inference on a toy problem. In particular, we build a flexible linear model based on a set of analytic basis functions that describe the simulated E-field beams of the HERA receiving elements with relatively few free parameters. An accurate and efficient basis that respects the symmetries of the problem is essential for avoiding significant model errors that could introduce spurious spectral and spatial structure. Achieving this in a compact way (i.e. with the fewest basis functions/coefficients possible) improves the prospect of directly inferring the structure of the beams from observations – particularly if this is to be done for each antenna within the array (rather than as an array average), out to the far sidelobes, and/or as a function of frequency. Conversely, a more ’cautious’, overly flexible basis would be harder to constrain with typical data, essentially only allowing us to marginalise over a very broad range of possible beam structures. The basis we propose in this paper has many of the properties required for a successful inference scheme in which individual antenna patterns can be measured versus frequency, and is tuneable in terms of its precision, at least with respect to the complex EM simulations that we use as a reference. Note that the methods we use in Paper I to determine this ‘sparse basis’ are not themselves Bayesian; the basis is in support of a Bayesian Gibbs sampling framework expounded in Paper II.
Analytic parametrization of the beam is an active area of research in the HERA collaboration (Choudhuri et al., 2021).111 http://reionization.org/manual_uploads/HERA101_Analytic_polarized_beam.pdf [^2] https://reionization.org/manual_uploads/HERA114_MEMO_beam_harmonics_Cynthia.pdf In general, analytic parametrizations have demonstrated significant promise in solving problems in HDR applications, e.g. more rapid beam evaluation (Asad et al., 2021), better physical beam characterization (Sekhar et al., 2022; Nasirudin et al., 2022), and removal of calibration biases that stem from gridding errors (Barry & Chokshi, 2022). This last reference also highlights the importance of accurate beam models for the sake of accurate gridding kernels in imaging power spectrum analyses (Morales et al., 2019).
We introduce a basis set that, to our knowledge, has not been used in the beam modeling literature. This basis solves several problems for us where other similar solutions from the literature did not.
  1. 1.
    It is fully analytic (unlike principle component analysis of holographic measurements for example). This makes it highly portable between different pixelization schemes while avoiding problems associated with discretization issues (e.g. spurious fluctuations in apparent source brightness at the horizon due to the setting of discrete pixel centers, or issues associated with sampling on sub-pixel scales for arrays with long baselines relative to the beam model resolution).
  2. 2.
    It describes the complex, realistic HERA beam simulations with relatively few basis functions compared to alternatives.
  3. 3.
    It behaves well at a coordinate singularity where e.g. polynomial-based bases did not (see §2).
  4. 4.
    The special functions involved have numerically stable scipy implementations up to high order which allows for thorough exploration of the parameter space.
We implement the sparse basis construction code of Paper I as well as the Gibbs sampling code of Paper II as a part of the Hydra Gibbs sampling framework (Kennedy et al., 2023). [^3] https://github.com/HydraRadio/Hydra The sparse fitting code makes use of the UVBeam class in the pyuvdata [^4] https://github.com/RadioAstronomySoftwareGroup/pyuvdata package (Hazelton et al., 2017). This class provides a general python interface for a variety of beam file formats, meaning the code we provide can be readily applied to other experiments.
@@ 2 Mathematical Formalism
In order to incorporate the beam into our Bayesian inference, we will need to posit a parameterized model for the beam that relates to the data from which the inference is drawn. In the general landscape of Bayesian inference, this is a rather arbitrary decision moderated only by the fact that some parametrizations may describe the data more efficiently than others. Without extreme prior knowledge about the beam, and in particular, uncertainties in the beam due to physical variations from the ideal scenario, we are left to choose some arbitrary basis for the beam with the hopes that our existing prior knowledge can at least constrain which basis functions are the most dominant. As a concrete example, we can take a detailed electromagnetic simulation or highly precise holographic measurement and fit a linear basis to the beam pattern for a high number of modes. Depending on the beam pattern and choice of basis functions, we may find that most of the fit is dominated by a small subset of basis functions. This has been shown to be relatively effective in de Lera Acedo et al. (2013); Young et al. (2013); Bui-Van et al. (2017); Iheanetu et al. (2019); Asad et al. (2021), though the choice of basis is often idiosyncratic to the telescope. We can then restrict ourselves to these dominant basis functions and consider the coefficients as the free parameters in our model.
@@
Figure 1: Jones components of the simulated Phase I E-W aligned dipole at 150 MHz, separated into real and imaginary components (rows), and by incident polarization (columns). The symbol $\mathcal{J}^{\mathrm{EW}}_{\mathrm{Az}}$ indicates the Jones component for the East-West aligned dipole in response to radiation polarized along the azimuth, while $\mathcal{J}^{\mathrm{EW}}_{\mathrm{ZA}}$ indicates the Jones component in response to radiation polarized along the zenith angle. There is an obvious dipolar structure, and higher azimuthal modes are clearly visible at zenith angles of 60 degrees or greater.
@@
An appealing alternative, though one that seems to us extremely computationally demanding for a full Bayesian computation, would be to have not only detailed knowledge of the ideal receiving element, but also have a parameterization for physical perturbations to this ideal along with detailed simulations or calculations of the response to them. Then a few parameters describing the perturbations only need to be constrained. See, for example, the characteristic basis function pattern approach in Maaskant et al. (2012); Young et al. (2013). The primary issue is that perturbations tend to belong to a multidimensional continuum and developing simulation outputs for finely graded perturbations to understand the resulting effect on the beam pattern is computationally prohibitive, though not infeasible depending on the study design (Kim et al., 2022, 2023). If a suitable emulator were constructed (i.e. if we could apply machine learning to produce fast but necessarily incomplete outputs of the simulator), the computational overhead could be reduced in the long term with an initial investment. However this might violate intellectual property law and is a somewhat ill-defined problem for a general electromagnetic simulator. Due to these seemingly intense challenges, and since our goal is mainly to establish the tractability and usefulness of our Gibbs sampling method, we opt for a more arbitrary linear basis expansion.
@@
Figure 2: Jones components of simulated HERA vivaldi feed at 150 MHz. Component notation and layout is identical to Figure 1. Spatial structure is somewhat more complex than the Phase I dipoles.
@@
We aim to infer the beam parameters from the visibilities via a forward model. We write the visibility equation in terms of the polarized beam response for antenna $j$ (Jones matrix, ignoring other propagation effects such as ionospheric refraction), $\boldsymbol{\mathcal{J}}_{j}(\hat{s},\nu)$, sky coherency matrix, $\boldsymbol{\mathcal{C}}(\hat{s},\nu)$, and physical antenna separation for antennas $j$ and $k$, $\Delta\vec{x}_{jk}$, as,
@@
where $c$ is the speed of light, $\nu$ is the frequency of observation, $\hat{s}$ is a unit vector pointing to different positions on the sky, $i^{2}=-1$, and depending on the coordinate system, some of these quantities are time-dependent. Different choices of coordinate system present different trade-offs. The Bayesian inference for the beam is drastically simpler if we operate in a coordinate system where the beam pattern is fixed. For the sky coordinates, we choose azimuth, $\phi$ and zenith angle, $\theta$. We choose sky polarization directions that align with this coordinate system. This puts all of the time dependence in the sky coherency matrix at the cost of producing a discontinuity in the beam pattern at zenith. This discontinuity arises because in this coordinate system, there is no well-defined direction for the polarization unit vectors at zenith. In math, if we think of the polarization unit vectors, $\hat{\theta}$ and $\hat{\phi}$ as vector fields, then
@@
In other words the projection is area-preserving (up to an overall factor of $2\alpha$). For $\alpha=1$, $\rho=0$ maps to zenith and $\rho=1$ maps to the horizon. The total volume contributed by a basis function to the square of a Jones element is simply given by
$V_{jp^{\prime}nm}^{p}=2\alpha|a_{jp^{\prime}nm}^{p}|^{2}.$ (8)
@@ 3 Compressing the Basis
There are two critical concerns when choosing a basis to represent the beam. First, assuming a priori knowledge of the beam through electromagnetic physics and knowledge of the receiving system, does the basis compactly represent the ideal beam assuming the perfect modeling of the receiving system? Second, assuming a class of perturbations to the receiving system, are ensuing perturbations to the beam pattern also compactly represented by the chosen basis? In this work, we provide a basis that satisfies the first question for the HERA receiving systems (both the Phase I dipole and Vivaldi feeds) using simulations of the ideal receiving system constructed in Fagnoni et al. (2021b, a), which use CST Microwave Studio (Weiland, 1977; Clemens & Weiland, 2001). We leave a full treatment of the second question for future work.
@@
In order to compress the basis, we need a way of determining which modes are the most important to use. Where appropriate data are available, various statistical methods may be used to perform this compression, such as principle component analysis of holographic measurements (Asad et al., 2021), or by inferring direction-dependent effects in the visibilities using a penalized likelihood (Yatawatta, 2018). We are opting to solve this compression problem before setting up the visibility based inference (in Paper II), rather than doing it on line during inference with e.g. a shrinkage estimator (Gelman et al., 2013). To do this, we examine equation 8 (the FBES) for the very precise fit using thousands of coefficients, shown in Figure 5 for both feed types and sky polarizations at 150 MHz. Almost all of the energy is at low $n$ and $m$ for both feeds, suggesting that we can choose a much smaller basis set with relatively little error. Since all of the plotted Jones elements have similar volume, we can interpret the number of high-energy modes as a proxy for the spatial complexity of the beam i.e. how compressible it is. For both feeds, we see that the response to azimuthally polarized radiation has longer tails to high radial modes compared to the response to zenith angle polarized radiation. Additionally, the Vivaldi feed has more higher energy modes than the Phase I dipole. For example the $|m|=3$ modes are prominent out to fairly high radial power, and there are more significant modes at much higher azimuthal number.
@@
Figure 10: (Top) Delay power spectra ($y$-axis) for the simulations described in section 4 using the HERA Dipole feed as a function of delay ($\tau$, $x$-axis) for the XX (left) and YY (right) polarizations. The solid black lines show the delay power spectra of the visibilities simulated with the CST beam from Fagnoni et al. (2021b). The colored lines show the delay power spectra for the simulations using our fits to the CST beam using various numbers of basis functions, $N_{\text{coeff}}$, as indicated in the colorbar. All power spectra have been normalized relative to the value of the black line at $\tau=0$. The grey shaded region marks the delays less than or equal to the Nyquist delay (500 ns) corresponding to the spectral resolution of the raw Fagnoni beam simulations (1 MHz). (Bottom) Fractional difference between the Fagnoni beam and FB simulations for each value of $N_{\text{coeff}}$. While the delay power spectra shown here correspond to a single 14.6 m EW baseline, the results are representative of all simulated baseline lengths and orientations.
@@
Figure 7 shows the RMSE as a function of zenith angle for both feed types and polarizations and a range of sparsity. We find that with 32 modes, there is $\sim 1\%$ agreement in the main lobe and $\sim 10\%$ agreement in the sidelobes. With 512 modes we achieve better than 1% precision at almost all zenith angles. The residuals from some of the sparse fits are shown in Figures 8 and 9 for the Phase I dipole and Vivaldi feeds, respectively. For a given level of sparsity, errors generally range by one order of magnitude above and below the RMSE value depending on the location of interest. We explore computational complexity for our Gibbs sampling pipeline in Paper II more thoroughly, however we expect having a beam model with 64-128 parameters is computationally feasible, while 512 modes will be tractable with a more optimized implementation.
We note that translating the observed errors in this fit to the actual errors produced by this approach in practice is not necessarily straightforward. The actual beam will at best be some perturbation of the simulated beam used in this work, not even including the close-packed array effects, which are known to vary even between different simulation packages on the order of a few percent (Bolli et al., 2023). Perhaps more importantly, the approach in Paper II involves inferring the beam based on the interferometric visibilities, whereas this least-squares fit is more analogous to a problem where one infers the beam based on a holographic map. In other words, this work verifies that the spatial structure of the beam can be captured by this basis with relatively few basis functions, but the errors shown do not necessarily represent the types of errors we may observe with a Bayesian point estimate since the information content of the visibilities will be different than this simulation. Furthermore, our approach will return a full Bayesian posterior, not just a point estimate. Folding uncertainties into a forward modeling procedure should hopefully alleviate some of the difficulties that come with point estimation such as calibration errors (Byrne et al., 2021).
@@ 4 Spectral Structure
We investigate any spurious spectral contamination introduced by fitting the beam with the chosen basis. To quantify this contamination, we compared the delay power spectra (Parsons & Backer, 2009; Parsons et al., 2012) from a suite of visibility simulations which use differing numbers of fit coefficients for the beam. All other input parameters to the simulations were kept fixed, however, to isolate the effects of changing the beam.
The delay spectrum is accessed by Fourier transforming (and then squaring) the visibilities along the frequency axis into “delay" space, so-named because the visibilities in this space represent timelike correlations between the antennas’ voltage signals. The intrinsic signal of astrophysical foregrounds is generally confined to delays less than the geometric delay of the baseline, but can be broadened due to instrumental effects (such as a chromatic beam or calibration errors), as well as the choice of spectral tapering function when performing the Fourier transform. Depending on baseline length and other chromatic instrumental effects, delays below a certain scale will be inaccessible without extremely accurate foreground subtraction, while delays greater than this will likely avoid foreground-related contamination (Datta et al., 2010; Morales et al., 2012; Parsons et al., 2012; Trott et al., 2012). For HERA, preserving high dynamic range for delays greater than $\sim$200-300 ns is essential to measurements of the 21-cm power spectrum during reionization (Parsons et al., 2012; HERA Collaboration et al., 2023). Therefore, the primary purpose of this test is to check whether there is a significant decrease in dynamic range above those delays.
@@
The simulations were performed using pyuvsim [^6] https://github.com/RadioAstronomySoftwareGroup/pyuvsim (Lanman et al., 2019), a high-precision visibility simulator which has been rigorously tested (Lanman et al., 2022; Aguirre et al., 2022). We simulated a selection of baselines from a close-packed hexagonal array layout with various orientations and lengths ($0\leq b\leq 90$ m). For the sky, we used the 2016 Global Sky Model (GSM, Zheng et al. (2017)), a map of galactic diffuse emission. For each antenna feed type (dipole or vivaldi), we ran simulations using varied numbers of beam fit coefficients, from 32 to 512 (“FB simulations”). For a given number of fit coefficients ($N_{\text{coeff}}$), we use the FBES to select the $N_{\text{coeff}}$ modes with the largest FBES amplitudes. We do this independently at each frequency, and take the union of all basis functions selected in this way across the simulated band. Since the top $N_{\text{coeff}}$ contributing basis functions varies slightly as a function of frequency, this means each selected $N_{\text{coeff}}$ is less than the number of modes actually used in each instance of the test. Note that this is different in previous sections, where we were only examining one frequency. As a reference against which we can compare the FB simulations, we also ran a simulation for each antenna feed type using the corresponding raw electromagnetic simulations from Fagnoni et al. (2021b, a) (“Fagnoni simulations”).
This test is greatly complicated by the fact that the reference beam simulations are only available at 1 MHz resolution, which cannot access most of the delays relevant to EoR science. To accommodate a more straight-forward comparison, we fit the Fourier-Bessel beam at 1 MHz resolution and then interpolated the coefficients (as well as the reference beam) during simulation using a cubic spline to a resolution of $\sim 97$ kHz to match the HERA spectral resolution. The Nyquist frequency for a 1 MHz bandwidth is 500 ns. Delays in the range $|\tau|>500$ ns are therefore susceptible to interpolation artifacts regardless of which beam is used. In other words, there is no obvious ground truth of what HERA should observe with this sky model above the Nyquist delay. With this caveat in mind, we list three important observations of the delay spectra from these visibility simulations, and then discuss them in more detail below. The delay spectra are shown in Figures 10 and 11 for the dipole and Vivaldi feeds, respectively.
  1. 1.
    The Fourier-Bessel beams generally produce slightly smoother visibilities than the simulated beams of the corresponding feed i.e. there is a loss of power at intermediate delays and beyond, meaning there is no hit to dynamic range.
  2. 2.
    There is a significant discrepancy below the Nyquist delay, visible in the XX delay spectra, which appear to have pronounced shoulders when using the reference beam. We argue that these are largely an artefact that stems from a complex interplay between spatial interpolation of the beam and the chosen sky model. If this is the case, this structure is spurious and it is desirable that the FB fits should not reproduce it.
  3. 3.
    There is also a significant discrepancy at $|\tau|\sim 1500$ ns in the XX beams, which is above the Nyquist delay. We suspect these are the pronounced shoulders from the reference beams being aliased to a harmonic of the Nyquist delay as a result of the frequency interpolation. This therefore represents high-delay contamination that is happily lacking when using the Fourier-Bessel beams.
@@
To better understand the choice of interpolation spline, we also ran a pyuvsim simulation using linear frequency interpolation of the Fagnoni beam. For linear interpolation, interpolation artifacts were highly visible as a set of peaks (duplicates of the peak at $|\tau|\leq 500$ ns) spaced uniformly every 500 ns (the Nyquist frequency for 1 MHz spectral resolution). While the amplitude of these duplicated peaks decreased with increasing $|\tau|$, the dynamic range between $\tau=0$ ns and the highest delays was only $\sim 10^{8}$. Whereas for the cubically interpolated Fagnoni simulations, we see a dynamic range of $\sim 10^{11}$ in Figures 10 and 11 and little evidence of such duplicated low-delay structure.
@@
There are, however, also differences between the Fagnoni beam and FB beam simulations inside the Nyquist delay. The raw Fagnoni beams were simulated on a rectilinear grid in altitude and azimuth with a 1 deg resolution. These beams thus also require spatial interpolation during visibility simulation to get the beam value at the location of each source (pixel) in the sky catalog. Our FB beam fits are analytic and thus interpolate smoothly to any pixelation scheme. To determine the importance of this spatial interpolation on our results, we also performed a set of simulations using a mock, gridded point source catalog with one point source at the center of each Fagnoni beam pixel at a single LST. In this case, the Fagnoni beams require no spatial interpolation. For this gridded point source catalog, we see almost no discrepancy inside the Nyquist delay. This suggests that the discrepancies we see for $|\tau|\sim 500$ ns are related to the spatial interpolation of the Fagnoni beam during visibility simulation i.e. it does not represent intrinsic spectral structure of the beam that is subsequently imprinted on any sky model for which it is used. However, since the gridded sky model does not have the same intrinsic spectral structure as the GSM, this additional test by itself leaves open the question as to whether these shoulders are intrinsic beam structure that somehow goes missing when we use the FB beams. By running the simulations of the GSM with an achromatic beam equal to the FB beam at the lowest interpolated frequency in the band, we find that these shoulders are not present, suggesting that the shoulders are not intrinsic to the GSM. In summary, we suspect the large discrepancies between the Fagnoni beam simulations and the best-fit FB beam (see Figures 10 and 11) are caused by interpolation issues with the simulated beams that the FB basis successfully avoids, rather than being due to a serious reconstruction error on the part of the FB basis.